# Automatic Hardware Pragma Insertion in High-Level Synthesis: A Non-Linear Programming Approach

### v2

@@ 7.5. Accuracy
The tightness of the lower bound estimation relies on the correct application of pragma directives such as pipeline and parallel. It also assumes that Merlin can efficiently transfer memory from off-chip to on-chip using 512-bit chunks. Finally, it assumes that Merlin optimally handles the transfer of memory from off-chip to on-chip. Figures 5a and 5b compare the measured HLS latency for every synthesizable design explored during our DSE with its predicted latency per solving the NLP. The Y-axis represents log(latency) and the X-axis the rank of the design sorted by HLS latency. For Fig 5b we exclude designs when we detect that the pragma parallel and pipeline are not applied as defined by Merlin. We observe that about half of the designs have at least one pragma not applied, leading logically to a larger difference between measured and predicted latency. Generally speaking, for parallelization pragmas, Merlin is more restrictive for coarse-grained parallelization, in many cases these pragmas are not applied. Coarse-grained pragmas are typically not applied to kernels that do not have an outermost reduction loop and thus can theoretically have coarse-grained parallelization, which is present in most linear algebra kernels such as 2mm, 3mm, gemver, etc. We also observe certain cases where the partitioning is not done correctly which does not allow a pipeline with II=1 when it is theoretically possible. For Figure 5b we observe a better overall accuracy, albeit imperfect. These differences are due in large part to how Merlin eventually implemented memory transfers, which we model optimistically. Internally, Merlin transforms the size of the arrays according to the program’s unroll factors and in certain cases does not allow transfers with a bitwidth of 512 bits.
(a) Comparison of the latency between the design reported in the HLS report and the lower bound estimate provided by the nonlinear problem for all explored designs
(b) Comparison of the latency between the design reported in the HLS report and the lower bound estimate provided by the nonlinear problem specifically for cases where pragmas were applied
Figure 5. Comparison of the latency between the design reported in the HLS report and the lower bound estimate provided by the nonlinear problem for all explored designs and specifically for cases where pragmas were applied.
@@
The results are presented in Table 6. On average, it takes 8 steps of the DSE to discover the design with the best QoR and 15 steps to terminate the DSE. We can observe that for some kernel we find the design with the best QoR at the first iteration of the DSE (which correspond to first shoot method in Table 5) but the DSE need more stop to stop the guarantee that we cannot obtain better latency, which implies that the lower bound is not perfectly tight.
@@ 8. Examples
In this section, we illustrate the significance of our method by contrasting it with AutoDSE and highlighting the advantages of our Design Space Exploration (DSE) approach. Our method excels in addressing domain-specific constraints, providing superior convergence in complex scenarios compared to AutoDSE. Furthermore, our DSE demonstrates adaptability and efficiency, proving to be robust in handling intricate design spaces, offering a more versatile and high-performing solution.
@@ 27 S3: D[i2][j2] += tmp[i2][k2] * C[k2][j2];
30 }
31 }
32}
Listing 11: Implementing the 2mm code with pragma directives for pipelining, tiling, and parallelization for each loop: $D=\alpha\times A\times B\times C+\beta\times D$.
2mm serves as a linear algebra kernel, acting as a surrogate for transformer inference, such as Bert. The code snippet in Listing 11 illustrates the Medium-sized configuration with potential pragma options. The PIPE pragma can be either flattened or off (default), while PARA and TILE can be any divisor of the loop trip count, defaulting to 1.
8.1.1. AutoDSE
  • three out of four of workers initially over-utilize parallelism by flattening PIPE_L0 and $\setminus$ or PIPE_L1 (and hence unroll the innermost loops), causing timeouts in current High-Level Synthesis (HLS) tools. Even without considering timeouts, these designs exceed array partitioning limits, preventing the reading of all data in a single cycle as expected by the unrolling process. Moreover, these designs strain hardware resources, requiring backtracking, which extends the search duration.
@@ 8.1.2. NLP-DSE
Now, we delve into how NLP-DSE overcomes these challenges to discover designs with superior QoR, emphasizing the usefulness of the Design Space Exploration (DSE) described in section 6.
The initial NLP-DSE design features parameters: PARA_L0 = 3, PIPE_L2 = flatten, PARA_L4 = 210, PARA_L1 = 6, PIPE_L3 = flatten, PARA_L5 = 190 achieving 13 GFLOPS/s. However, the compiler fails to apply PARA_L0 and PARA_L1 pragmas, creating a performance gap.
The second design is with the parameters: PIPE_L2 = flatten, PARA_L2 = 2, PARA_L4 = 210, PARA_L3 = 5, PIPE_L3 = flatten, PARA_L5 = 190 which allows to achieve a throughput of 85 GFLOPS/s. However the unroll factor of PARA_L3 does not allow the compiler Xilinx Merlin to transfer the data from off-chip to on-chip efficiently which does not allow to achieve the lower bound. Hence we continue the search and we found our the design with the best QoR in our space at the iteration 8 with the parameters: PIPE_L2 = flatten, PARA_L2 = 2, PARA_L4 = 210, PARA_L3 = 2, PIPE_L3 = flatten, PARA_L5 = 190 We can note that Xilinx Merlin did not allow the data to be transferred optimally so we did not achieve the lower bound and we continue the search. From iteration 10 the lower bound found by the NLP is greater than the latency (HLS report) of design 8, which makes it possible to stop the search because even if we reach the lower bound we will have a latency greater than what we have already obtained.
The Figure 6 summarize the result achieved at each step of the DSE. The step 4,5 and 6 (red) found the same configurations as the step 2 and 3 so we do not need to run the synthesis as we already have the results.
@@ 9. Related Work
NLP-DSE makes it possible to automatically introduce pragmas in order to obtain a design with a good QoR. Many previous works using different DSE methods have the same objective. These model-free Design Space Exploration (DSE) techniques, as exemplified by works such as (Sohrabizadeh et al., 2021; Yu et al., 2018; Ferretti et al., 2018), employ a methodology where the compiler acts as a black-box, and they dynamically adapt their exploration strategies based on the outcomes of previous iterations. In these approaches, each candidate design is evaluated by generating a High-Level Synthesis (HLS) report. However, the time required for the synthesis or report generation can extend over several hours, significantly limiting the breadth of the explored design space. Moreover, it is worth noting that certain DSE methodologies, including those described in (Sohrabizadeh et al., 2021), might encounter challenges such as converging to local minima, which can impede the discovery of the globally optimal solution.
@@
To circumvent the constraints imposed by synthesis time, novel approaches in Design Space Exploration (DSE) have emerged, including model-based DSEs and AI-driven DSEs. These methods leverage sophisticated techniques such as cost modeling (Zhao et al., 2017; Bannwart Perina et al., 2019; Zhong et al., 2016; Zuo et al., 2015a), Neural Networks (NN) (Wang et al., 2017; Zhong et al., 2017; Koeplinger et al., 2016; Liu and Carloni, 2013; Liu et al., 2019; Schafer and Wakabayashi, 2012; Xydis et al., 2015), Graph Neural Networks (GNN) (Sohrabizadeh et al., 2023b, 2022; Wu et al., 2023), or decision trees (DT) (O’Neal et al., 2018; Yu et al., 2021; Makrani et al., 2019) to estimate the Quality of Results (QoR) of each design rapidly. By utilizing these techniques, the evaluation time for a single design can be reduced to mere milliseconds. However, despite this acceleration, assessing a large number of designs still entails a significant time investment. Furthermore, while these rapid evaluations provide valuable insights, they may not perfectly align with the outcomes obtained from High-Level Synthesis (HLS) reports in terms of accuracy. Consequently, relying solely on HLS validation for the top-$n$ results may lead to suboptimal solutions, as the rapid evaluation methods might not capture all pertinent design intricacies. Therefore, a more comprehensive approach that combines the strengths of both rapid evaluation techniques and traditional HLS validation is necessary to ensure optimal design outcomes.
@@
In contrast, alternative methodologies offer one-shot optimization through code transformations and pragma insertion, as evidenced by works like (Ye et al., 2022; Lai et al., 2019; Huang et al., 2021). However, the efficacy of these approaches is constrained by the limited scope of available hardware directives and code transformations. While some endeavors concentrate on predefined micro-architectures, as seen in (Wang et al., 2021; Basalama et al., 2023b), their applicability is restricted. Moreover, specialized applications such as Deep Neural Networks (DNN) (Zhang et al., 2018, 2020), stencil computations (Chi et al., 2018), sparse linear algebra operations (Du et al., 2022), and neural networks (Basalama et al., 2023a; Sohrabizadeh et al., 2020), including Convolutional Neural Networks (CNNs) (Reggiani et al., 2019), have garnered attention. Yet, these methods encounter challenges when extrapolated beyond their designated domains, rendering generalizations difficult. Hence, while these approaches offer streamlined optimization strategies and tailored solutions for specific problem domains, their broader applicability beyond their respective niches remains a challenge.
@@
NLP-DSE presents a hybrid methodology, leveraging a NLP-based cost model to swiftly explore expansive design spaces within minutes, potentially outpacing existing models in speed. Nevertheless, to ensure precise performance evaluation, reliance on High-Level Synthesis (HLS) remains integral to our approach. Recent advances in optimization solvers such as BARON (Tawarmalani and Sahinidis, 2005; Sahinidis, 2017) have allowed NLP-based approaches to become a promising alternative to approximate ILP-based methods, as they can encode more complex and realistic performance models. Unlike prior works (Bondhugula et al., 2008; Pouchet et al., 2013b; Zuo et al., 2017) that frame the cost model as Linear Programming (LP) problems, NLP-based methods can handle more complex constraints without necessitating approximations, such as estimating communication volumes across loops (Bondhugula et al., 2008; Pouchet et al., 2013b) or simplifying the space by exposing direct parallelization in the problem (Zuo et al., 2017). It is noteworthy that the comparison landscape lacked other NLP-based methodologies, while Linear Programming methods were deemed less suitable due to their incapacity to manage nonlinear constraints effectively. The inclusion of multiple product terms within the objective function and constraints, such as the product of Unroll Factors (UF) for perfectly nested loops, underscores the necessity for accurate modeling, which is challenging to approximate with linear approximations.
@@
Approaches that do not rely on precise static analysis, such as (Sohrabizadeh et al., 2021; Yu et al., 2018; Ferretti et al., 2018; Sohrabizadeh et al., 2022; Wu et al., 2023), can take as input any C/C++ kernel supported by the HLS compiler, thus expanding their applicability to a broader spectrum of programs. In contrast, NLP-DSE , akin to other model-based Design Space Exploration (DSE) approaches (Zhao et al., 2017; Bannwart Perina et al., 2019; Zhong et al., 2016; Zuo et al., 2015a), is constrained to affine programs to ensure precise analysis and facilitate the modeling of latency and resource utilization. While this encompasses a significant subset of programs, including AI kernels, and aligns with the MLIR affine dialect, it does not match the versatility of other frameworks. Consequently, there exists a tradeoff between accuracy and generality. Opting for a subset of programs provides more detailed information, thereby accelerating DSE and potentially improving the QoR.

### v4

@@ 8.6. Accuracy
The tightness of the lower bound estimation relies on the correct application of pragma directives such as pipeline and parallel. It also assumes that Merlin can efficiently transfer memory from off-chip to on-chip using 512-bit chunks. Finally, it assumes that Merlin optimally handles the transfer of memory from off-chip to on-chip. Figures 4a and 4b provide a comprehensive comparison between the measured HLS latency for every synthesizable design explored during our DSE process and the predicted latency obtained by solving the NLP.
The Y-axis represents log(latency) and the X-axis the rank of the design sorted by HLS latency. For Fig 4b we exclude designs when we detect that the pragma parallel and pipeline are not applied as defined by Merlin. We observe that about half of the designs have at least one pragma not applied, leading logically to a larger difference between measured and predicted latency. Generally speaking, for parallelization pragmas, Merlin is more restrictive for coarse-grained parallelization, in many cases these pragmas are not applied. Coarse-grained pragmas are typically not applied to kernels that do not have an outermost reduction loop and thus can theoretically have coarse-grained parallelization, which is present in most linear algebra kernels such as 2mm, 3mm, gemver, etc. We also observe certain cases where the partitioning is not done correctly which does not allow a pipeline with II=1 when it is theoretically possible. For Figure 4b we observe a better overall accuracy, albeit imperfect. These differences are due in large part to how Merlin eventually implemented memory transfers, which we model optimistically. Internally, Merlin transforms the size of the arrays according to the program’s unroll factors and in certain cases does not allow transfers with a bitwidth of 512 bits.
@@
(a) For all explored designs
(b) For cases where pragmas were applied as expected
Figure 4. Comparison of the Latency Between the Design Reported in the HLS Report and the Lower Bound Estimate Provided by the Nonlinear Problem for All Explored Designs, Specifically for Cases Where Pragmas Were Applied as Expected, with ID Rank Representing the Order of Designs Sorted by HLS Latency.
@@
The results are presented in Table 8. On average, it takes 8 steps of the DSE to discover the design with the best QoR and 15 steps to terminate the DSE. We can observe that for some kernel we find the design with the best QoR at the first iteration of the DSE (which correspond to first shoot method in Table 6) but the DSE needs more step to stop the guarantee that we cannot obtain better latency, which implies that the lower bound is not perfectly tight.
@@ 9. Examples
In this section, we illustrate the significance of our method by contrasting it with AutoDSE and highlighting the advantages of our DSE approach. Our method excels in addressing domain-specific constraints, providing superior convergence in complex scenarios compared to AutoDSE. Furthermore, our DSE demonstrates adaptability and efficiency, proving to be robust in handling intricate design spaces, offering a more versatile and high-performing solution.
@@
2mm serves as a linear algebra kernel, acting as a surrogate for transformer inference, such as Bert. The code snippet in Listing 6 illustrates the Medium-sized configuration with potential pragma options. The PIPE pragma can be either flattened or off (default), while PARA and TILE can be any divisor of the loop trip count, defaulting to 1.
9.1.1. AutoDSE
  • three out of four of workers initially over-utilize parallelism by flattening PIPE_L0 and/or PIPE_L1 (and hence unroll the innermost loops), causing timeouts in current High-Level Synthesis (HLS) tools. Even without considering timeouts, these designs exceed array partitioning limits, preventing the reading of all data in a single cycle as expected by the unrolling process. Moreover, these designs strain hardware resources, requiring backtracking, which extends the search duration.
@@ 27 S3: D[i2][j2] += tmp[i2][k2] * C[k2][j2];
30}
Listing 6: Implementing the 2mm code with pragma directives for pipelining, tiling, and parallelization for each loop: $D=\alpha\times A\times B\times C+\beta\times D$.
@@ 9.1.2. NLP-DSE
Now, we delve into how NLP-DSE overcomes these challenges to discover designs with superior QoR, emphasizing the usefulness of the DSE described in section 7.
The initial NLP-DSE design features parameters (Step 1 in Figure 5): PARA_L0 = 3, PIPE_L2 = flatten, PARA_L4 = 210, PARA_L1 = 6, PIPE_L3 = flatten, PARA_L5 = 190 achieving 13 GFLOPS/s. However, the compiler fails to apply PARA_L0 and PARA_L1 pragmas, creating a performance gap.
The second design (Step 2 in Figure 5) uses the following parameters: PIPE_L2 = flatten, PARA_L2 = 2, PARA_L4 = 210, PARA_L3 = 5, PIPE_L3 = flatten, and PARA_L5 = 190. This configuration achieves a throughput of 85 GFLOPS/s. However, the unroll factor for PARA_L3 does not allow the Xilinx Merlin compiler to efficiently transfer data from off-chip to on-chip, preventing the design from reaching the lower bound performance.
In Step 3, the design parameters are: PARA_L0 = 3, PIPE_L2 = flatten, PARA_L4 = 210, PARA_L1 = 3, PARA_L3 = 2, PIPE_L3 = flatten, and PARA_L5 = 190. This configuration achieves a throughput of 65 GFLOPS/s, which is lower than the design found in Step 2, as the compiler did not apply the PARA_L0 pragma as expected. Steps 4, 5, and 6 (marked in red) identify the same configurations as Steps 2 and 3, so synthesis is not necessary as the results are already available. For Step 7, the HLS compiler is unable to apply the configuration (PARA_L0 = 2, PIPE_L2 = flatten, PARA_L4 = 210, PARA_L1 = 6, PARA_L3 = 44, PIPE_L5 = flatten, PARA_L5 = 5), leading to a throughput of only 0.82 GFLOPS/s.
The design with the best QoR in our search space is found at Step 8 of the DSE, with parameters: PIPE_L2 = flatten, PARA_L2 = 2, PARA_L4 = 210, PARA_L3 = 2, PIPE_L3 = flatten, and PARA_L5 = 190. However, due to Xilinx Merlin’s inability to optimally transfer data, the lower bound is not achieved, and the search continues.
In iteration 9, a design is found with the parameters PIPE_L2 = flatten, PARA_L4 = 210, PARA_L1 = 6, PARA_L3 = 22, PIPE_L5 = flatten, and PARA_L5 = 10. The pragma are not applied as expected, and the throughput is only 1.35 GFLOPS/s. From iteration 10 the lower bound found by the NLP is greater than the latency (HLS report) of design 8, which makes it possible to stop the search because even if we reach the lower bound we will have a latency greater than what we have already obtained. The Figure 5 summarize the result achieved at each step of the DSE.
@@ 10. Related Work
NLP-DSE makes it possible to automatically introduce pragmas in order to obtain a design with a good QoR. Many previous works using different DSE methods have the same objective. These model-free Design Space Exploration (DSE) techniques, as exemplified by works such as (Sohrabizadeh et al., 2022b; Yu et al., 2018; Ferretti et al., 2018), employ a methodology where the compiler acts as a black-box, and they dynamically adapt their exploration strategies based on the outcomes of previous iterations. In these approaches, each candidate design is evaluated by generating a High-Level Synthesis (HLS) report. However, the time required for the synthesis or report generation can extend over several hours, significantly limiting the breadth of the explored design space. Moreover, it is worth noting that certain DSE methodologies, including those described in (Sohrabizadeh et al., 2022b), might encounter challenges such as converging to local minima, which can impede the discovery of the globally optimal solution.
@@
To circumvent the constraints imposed by synthesis time, novel approaches in Design Space Exploration (DSE) have emerged, including model-based DSEs and AI-driven DSEs. These methods leverage sophisticated techniques such as cost modeling (Zhao et al., 2017; Bannwart Perina et al., 2019; Zhong et al., 2016; Zuo et al., 2015a), Neural Networks (NN) (Wang et al., 2017; Zhong et al., 2017; Koeplinger et al., 2016; Liu and Carloni, 2013; Liu et al., 2019; Schafer and Wakabayashi, 2012; Xydis et al., 2015), Graph Neural Networks (GNN) (Sohrabizadeh et al., 2023b, 2022a; Wu et al., 2023), or decision trees (DT) (O’Neal et al., 2018; Yu et al., 2021; Makrani et al., 2019) to estimate the Quality of Results (QoR) of each design rapidly. By utilizing these techniques, the evaluation time for a single design can be reduced to mere milliseconds. However, despite this acceleration, assessing a large number of designs still entails a significant time investment. Furthermore, while these rapid evaluations provide valuable insights, they may not perfectly align with the outcomes obtained from High-Level Synthesis (HLS) reports in terms of accuracy. Consequently, relying solely on HLS validation for the top-$n$ results may lead to suboptimal solutions, as the rapid evaluation methods might not capture all pertinent design intricacies. Therefore, a more comprehensive approach that combines the strengths of both rapid evaluation techniques and traditional HLS validation is necessary to ensure optimal design outcomes.
In contrast, alternative methodologies offer one-shot optimization through code transformations and pragma insertion, as evidenced by works like (Ye et al., 2022; Lai et al., 2019; Huang et al., 2021). However, the efficacy of these approaches is constrained by the limited scope of available hardware directives and code transformations. While some endeavors concentrate on predefined micro-architectures, as seen in (Wang et al., 2021; Basalama et al., 2023b), their applicability is restricted. Although ScaleHLS offers valuable code transformations that are not implemented in our approach, our framework provides a more comprehensive design space for pragma insertion, which is exhaustively explored within seconds or minutes using our NLP solver. While ScaleHLS uses a QoR estimator for each pragma configuration—despite its quick estimation times (e.g., milliseconds)—it cannot explore large design spaces exhaustively. This limitation is similar to those observed with HARP. In contrast to ScaleHLS, our design-space exploration (DSE) approach focuses on more than just rapidly identifying a theoretical solution. Our NLP solver thoroughly explores the entire design space in seconds or minutes to find the optimal theoretical solution. Additionally, our DSE approach ensures that pragmas are correctly applied by the compiler through synthesis, addressing a gap that ScaleHLS does not cover. Furthermore, ScaleHLS employs the Vivado flow (Xilinx, 2024), which assumes that data are on-chip. Thus, they consider a less complex space compared to ours, as we also account for the memory transfer from off-chip to on-chip.
@@
Moreover, specialized applications such as Deep Neural Networks (DNN) (Zhang et al., 2018, 2020), stencil computations (Chi et al., 2018), sparse linear algebra operations (Du et al., 2022), and neural networks (Basalama et al., 2023a; Sohrabizadeh et al., 2020), including Convolutional Neural Networks (CNNs) (Reggiani et al., 2019; Mousouliotis et al., 2023), have garnered attention. Yet, these methods encounter challenges when extrapolated beyond their designated domains, rendering generalizations difficult. Hence, while these approaches offer streamlined optimization strategies and tailored solutions for specific problem domains, their broader applicability beyond their respective niches remains a challenge.
@@
NLP-DSE presents a hybrid methodology, leveraging a NLP-based cost model to swiftly explore expansive design spaces within minutes, potentially outpacing existing models in speed. Nevertheless, to ensure precise performance evaluation, reliance on High-Level Synthesis (HLS) remains integral to our approach. Recent advances in optimization solvers such as BARON (Tawarmalani and Sahinidis, 2005; Sahinidis, 2017) have allowed NLP-based approaches to become a promising alternative to approximate ILP-based methods, as they can encode more complex and realistic performance models. Unlike prior works (Bondhugula et al., 2008; Pouchet et al., 2013; Zuo et al., 2017) that frame the cost model as Linear Programming (LP) problems, NLP-based methods can handle more complex constraints without necessitating approximations, such as estimating communication volumes across loops (Bondhugula et al., 2008; Pouchet et al., 2013) or simplifying the space by exposing direct parallelization in the problem (Zuo et al., 2017). It is noteworthy that the comparison landscape lacked other NLP-based methodologies, while Linear Programming methods were deemed less suitable due to their incapacity to manage nonlinear constraints effectively. The inclusion of multiple product terms within the objective function and constraints, such as the product of Unroll Factors (UF) for perfectly nested loops, underscores the necessity for accurate modeling, which is challenging to approximate with linear approximations.
@@
Approaches that do not rely on precise static analysis, such as (Sohrabizadeh et al., 2022b; Yu et al., 2018; Ferretti et al., 2018; Sohrabizadeh et al., 2022a; Wu et al., 2023, 2021), can take as input any C/C++ kernel supported by the HLS compiler, thus expanding their applicability to a broader spectrum of programs. In contrast, NLP-DSE, akin to other model-based Design Space Exploration (DSE) approaches (Zhao et al., 2017; Bannwart Perina et al., 2019; Zhong et al., 2016; Zuo et al., 2015a), is constrained to affine programs to ensure precise analysis and facilitate the modeling of latency and resource utilization. While this encompasses a significant subset of programs, including AI kernels, and aligns with the MLIR affine dialect, it does not match the versatility of other frameworks. Consequently, there exists a tradeoff between accuracy and generality. Opting for a subset of programs provides more detailed information, thereby accelerating DSE and potentially improving the QoR.